# Matching Markets Meet LLMs: Algorithmic Reasoning with Ranked Preferences

### v1

@@ Acknowledgments
This research was supported in part by NSF Awards IIS-2144413 and IIS-2107173.
@@ Appendix A Limitations and Future Work
Additionally, while our work provides insights into the reasons behind the failure of LLMs to consistently generate stable solutions (see Section 6), there is scope for further clarity on where exactly LLMs make mistakes during algorithmic execution. A potential method to explore this is to break the algorithmic execution task into smaller steps (e.g. a single proposal-acceptance/rejection cycle) and identify which components of the state-transition are challenging for LLMs to understand.
Furthermore, while fine-tuning substantially enhances LLMs’ performance on instances with relatively smaller input sizes, improving their performance with larger inputs requires further exploration. This can include the investigation of methods such as fine-tuning the entire set of parameters (unlike with LoRA) or reinforcement-learning methods such as group-relative policy optimization (GRPO) that are known to increase the reasoning capabilities of LLMs [61].
@@ C.1 Algorithm for Generating a Stable Matching
As shown in Algorithm 1, the standard deferred-acceptance algorithm runs by having the proposing side of the market make a series of proposals, and each agent that receives at least one proposal decides which proposal to accept (the proposal becomes an engagement), and which proposals to reject (or engagements to break). This continues until all agents are matched, which requires $O(n^{2})$ proposal steps. The resulting solution is stable [27]. To describe the algorithm, we adopt the traditionally used setting of stable-marriage where the proposing side consists of men and the receiving side consist of women.
@@
Another commonly used version of the DA algorithm involves the receiver agent (in a given proposal) removing all agents (on the proposer-side) from their preference list who are ranked below the current proposer agent (who also remove the receiver from their respective preference lists). Due to the shortening of the preference lists as the algorithm progresses (see Algorithm 2), this is referred to as the version of the DA algorithm with Shortlists [36]. While this version terminates with at most as many (and often less) proposal steps as compared to Algorithm 1, it requires repeated updates to the original preference lists.
@@
C.2 Algorithm for Generating a Stable Matching w/ Master Lists
Similar to Algorithm 1, Algorithm 3 runs in rounds of proposals. Since there is a Master-list over proposing agents in the preference lists of receiving agents, and proposing agents are selected to make proposals in the order in which they appear in the Master-list, there are no rejections (since any receiver receives the best possible proposal at any step) [37]. Hence, this algorithm terminates in $n$ proposal steps (and is therefore easier to execute).
Algorithm 3 The Deferred-Acceptance Algorithm for Preferences w/ Master Lists on One Side
@@ C.3 Algorithm for Resolving Instability
While we don’t explicitly describe the algorithm here, the mechanism presented by Abeledo and Rothblum [1] can be applied to an unstable matching $\mu$ by resolving blocking pairs, resulting in a stable solution. Intuitively, an LLM does not have to follow a specific mechanism, rather the model can resolve instability by iteratively resolving blocking pairs as they arise (eventually, assuming all steps are correct, the model should arrive at a stable solution).
@@ C.4 Algorithm for Detecting Instability
Intuitively, Algorithm 4 works by iteratively visiting each pair of agents $(m,w)$ s.t. $m\in M$ and $w\in W$, and finding a pair such that either $m$ prefers $w$ to their current partner, or $w$ prefers $m$ to their current partner (when such a pair is found, output it as a blocking pair). If no pair $(m,w)$ is found to be a blocking pair, then the solution is stable.
@@ Appendix D Prompt Engineering
In Table 4, examine LLMs’ performance when provided with a prompt containing pseudocode for the DA algorithm, and compare it to the case when no algorithm is provided in the prompt. While providing the DA algorithm as part of the prompt leads to an improvement in some cases, the only case in which there is improvement is statistically significant [^9] At $p<0.05$, using Fisher’s exact test is with DeepSeek-70B at the Easy level with ML instances.
@@
For models that fail at generating stable matchings even with small instances, we evaluate whether prompt-based enhancements such as Chain-of-Thought (CoT) [70] and Few-shot prompting [19] can improve their performance. In particular, we introduce the following three types of modifications to the original prompt (used in Section 3):
  • CoT-Vanilla (CoT-V): The steps of execution of the DA algorithm (see Algorithm 1) are provided for an example instance. Each step consists of a single (free) proposing agent making his next proposal, and the receiving agent either accepting or rejecting the proposal based on their current status.
  • CoT-Shortlist (CoT-SL): This version of CoT uses the Shortlist version of the DA algorithm (see Algorithm 2) which reduces the overall number of proposal steps, but requires repeated updates to the original preference lists.
@@ 20 IC 0 0 0 0 0 0 0 0 14 2 6 6 0 0 0 0
As depicted in Table 5, these prompting enhancements fail to improve the ability of LLMs to generate stable matchings. While there is a slight improvement for models like Qwen-QwQ-32B and DeepSeek-70B instances with Master-list preferences and size $n=10$, this improvement is not statistically significant.
@@
We consider the traditionally used setting of stable-marriage, where the set $M$ consists of men who propose to women in the set $W$ [26]. To measure whether LLMs are sensitive to the nomenclature used to described the two-sided matching market, we also consider a different setting, i.e. that where a set of workers ($W$) are to be assigned a set of tasks ($T$) (and members on both sides have preferences over members of the other). We test the difference in the performance of two LLMs—Gemini-2.0-Flash and o3-mini—between the task-scheduling and stable-marriage settings. The results are provided in Table 6. While there is a slight decrease in performance for o3-mini, the change is not significant (at $p<0.05$). This provides further evidence that LLMs understand requirements of computing stable solutions for matching markets, in general.
@@ 50 IC 0 0 80 72
Appendix E Resolving Instability: Additional Material
E.1 Generating Unstable Matchings
@@
We generate a matching that contains a single blocking pair, by swapping the partners of two randomly selected proposer agents in the Optimal matching. Since such a swap may lead to more than one blocking-pair (or no blocking pairs), we perform this process (for every instance) until we obtain a matching with exactly one blocking-pair. This procedure is formally described in Algorithm 5.
@@
2:$\,\Pi=(\succ_{m},\succ_{w})$triangleright Preference profile for all men $m\in M$ and women $w\in W$
3:$\mu^{\ast}$triangleright Men-optimal stable matching returned by Deferred Acceptance
@@
7:     $\triangleright$ Keep trying until the condition is met
8:         $\mu\leftarrow\text{copy}(\mu^{\ast})$triangleright Start from the stable matching
@@
11:         $w_{b}\leftarrow\mu(m_{b})$triangleright Swap partners of the two men
@@
E.2 Further Results
Table 7: Instability Rate (averaged across instances) in the (valid) matchings returned by LLMs when asked to resolve a given unstable matching of types One-BP and Random. The column “Baseline” contains the (average) Instability Rate for the provided matching of the indicated type. Numbers in bold indicate that the Instability Rate of the corrected solution is significantly worse than the provided matching. A * on the number in the One-BP column indicates a that Instability Rate is significantly lower than the case when a Random matching is provided (at $p<0.05$). [^13] All statistical comparisons in this table are made using Welch’s t-test [71].
@@ E.2 Further Results
The extent to which a given matching is incorrect does influence quality of the solutions returned by LLMs after being asked to correct it. Table 7 shows that the number of blocking pairs is often significantly lower when the unstable matching (that LLMs are asked to correct) has a single blocking pair, as compared to the case with random matchings. [^14] In fact, asking LLMs to resolve a random matching leads to a significantly higher number of blocking pairs in the returned solution, as compared to the case when they are asked to generate solutions from scratch. Similarly, as shown in Figure 6, matchings returned after resolving an almost stable matching have a greater overlap with the Optimal solution (especially at Medium and Hard difficulty levels).
Appendix F Additional Material about Detecting Instability
F.1 Comparing Impartial Culture and Master-List Instances
@@
Appendix G Fine-tuning Details
@@
We fine-tuned three reasoning models:
@@
  • DeepSeek-14B (deepseek-ai/DeepSeek-R1-Distill-Qwen-14B), and
  • Qwen-QwQ-32B (Qwen/QwQ-32B),
using the Unsloth [^15] https://unsloth.ai/ framework with parameter-efficient tuning (LoRA).
@@
The dataset for the Generation task contains $N=10000$ samples for which the reasoning trace is generated using a Python implementation of the DA algorithm. For the Preference Reasoning task, the dataset consists of $N=9000$ samples ($3000$ for each question level), where the reasoning trace involves explicitly identifying the positions of agents in the concerned preferences. In both datasets, we include an equal number of IC and ML instances, with sizes ranging from $n=5$ to $n=50$. Detailed examples of training examples for both tasks (Generation and Preference Reasoning) are provided in Appendix J.
@@
Appendix H Inference Details
@@
Appendix I Prompts
I.1 Example prompt for Generating Stable Solutions
@@
The prompt used for generating stable solutions with LLMs follows standard prompting procedures, by first outlining the task, providing appropriate context, specifying constraints, and detailing the desired output format (a JSON object). Notice that despite the deferred-acceptance algorithm never being mentioned in the prompt, all models mentioned the deferred-acceptance algorithm in their responses. As mentioned in Appendix D, we use the traditional setting of stable-marriage (where men propose to women) considered by Gale and Shapley [27] while describing the problem in the prompt.
You are an intelligent assistant who is an expert in algorithms. Consider the following instance of the two-sided matching problem, where 10 men are to be matched with 10 women. Here are the preference lists for all individuals: <<<preferences>>> { M: { M1: [W10,W1,W3,W6,W2,W4,W9,W8,W7,W5], M2: [W8,W3,W10,W6,W2,W5,W4,W7,W1,W9], M10: [W2,W5,W1,W3,W7,W6,W10,W4,W9,W8], }, W: { W1: [M2,M8,M9,M10,M5,M7,M1,M4,M6,M3], W2: [M2,M7,M3,M1,M8,M9,M6,M10,M5,M4], W10: [M6,M4,M7,M5,M8,M9,M10,M2,M3,M1], }} <<</preferences>>> Your task is to find the proposer-optimal stable matching. You can use XML tags like <<<scratchpad>>> to explain your thought process while computing the solution. Once you have found a stable matching, please return your matching in the JSON format given below: <<<answer>>> { “M1”: “<<<woman matched with M1>>>”, “M2”: “<<<woman matched with M2>>>”, “M10”: “<<<woman matched with M10>>> } <<</answer>>> Make sure that each man/woman is matched with exactly ONE partner. It is mandatory that you provide a matching as a JSON object enclosed in <<<answer>>><<</answer>>> tags as described above.
@@
You are an intelligent assistant who is an expert in algorithms. $<$/preferences$>$ Your task is to find the proposer-optimal stable matching. For this, you can use the Deferred Acceptance algorithm. The steps of this algorithm are described below: 1. Initialize all men and women as unmatched. 2. Create a list to keep track of each man’s next proposal (initially set to 0 for all men). 3. While there are unmatched men: a. Select an unmatched man (M). b. Find the next woman (W) on M’s preference list that he hasn’t proposed to yet. c. If W is unmatched, match M and W. d. If W is matched but prefers M to her current partner: - Unmatch W from her current partner. - Match M and W. - Set the unmatched man as W’s previous partner. e. If W rejects M, move to the next woman on M’s preference list. 4. Repeat step 3 until all men are matched. You can use XML tags like $<$scratchpad$>$ to explain your thought process … It is mandatory that you provide a matching as a JSON object enclosed in $<$answer>>><<</answer$>$ tags as described above.
@@
You are an intelligent assistant who is an expert in algorithms. Consider the following instance of the two-sided matching problem, where 5 workers are to be assigned with 5 tasks, and each worker is assigned exactly one task. Here are the preference lists for all workers (W) over tasks (T) and the preferences of tasks over workers: <<<preferences>>> { W: { W1: [T5, T3, T4, T2, T1] } T: { T1: [W3, W5, W4, W1, W2] }} <<</preferences>>> Your task is to find a stable matching of workers and tasks. You can use XML tags like <<<scratchpad>>> to explain your thought process while computing the solution. Once you have found a stable matching, please return your matching in the JSON format given below: <<<answer>>> { "W1": "<<<task assigned to W1>>>", "W5": "<<<task assigned to W5>>>" } <<</answer>>> Make sure that each worker is assigned exactly ONE task. It is mandatory that you provide a matching as a JSON object enclosed in <<<answer>>><<</answer>>> tags as described above.
I.2 Example Prompts for Prompt Engineering
I.2.1 CoT-Vanilla
@@
Chain-of-Thought methods were applied to the prompt in Section I.1 by additionally including an example trace of steps performed when running the deferred-acceptance algorithm on a randomly generated instance. The algorithm trace includes all proposals, all respective acceptances/rejections, and the resultant stable solution. The entire Chain-of-Thought example is enclosed within $<$example$>$ XML tags.
[…]
I.2.2 CoT-Shortlist

### v2

@@ Acknowledgments
This research was supported by the National Science Foundation (NSF) through CAREER Award IIS-2144413 and Award IIS-2107173. We thank the anonymous reviewers for their fruitful comments. We would also like to thank Shraddha Pathak for several useful discussions.
@@ Appendix A Limitations and Future Work
Additionally, while our work provides insights into the reasons behind the failure of LLMs to consistently generate stable solutions (see Section˜6), there is scope for further clarity on where exactly LLMs make mistakes during algorithmic execution. A potential method to explore this is to break the algorithmic execution task into smaller steps (e.g., a single proposal-acceptance/rejection cycle) and identify which components of the state-transition are challenging for LLMs to understand.
Furthermore, while fine-tuning substantially enhances LLMs’ performance on instances with relatively smaller input sizes, improving their performance with larger inputs requires further exploration. This can include the investigation of methods such as fine-tuning the entire set of parameters (unlike with LoRA) or reinforcement-learning methods such as group-relative policy optimization (GRPO) that are known to increase the reasoning capabilities of LLMs [shao2024deepseekmathpushinglimitsmathematical].
@@ C.1 Algorithm for Generating a Stable Matching
As shown in Algorithm˜1, the standard deferred-acceptance algorithm runs by having the proposing side of the market make a series of proposals, and each agent that receives at least one proposal decides which proposal to accept (the proposal becomes an engagement), and which proposals to reject (or engagements to break). This continues until all agents are matched, which requires $O(n^{2})$ proposal steps. The resulting solution is stable [gale1962college]. To describe the algorithm, we adopt the traditionally used setting of stable-marriage where the proposing side consists of men and the receiving side consist of women.
@@
Another commonly used version of the DA algorithm involves the receiver agent (in a given proposal) removing all agents (on the proposer-side) from their preference list who are ranked below the current proposer agent (who also remove the receiver from their respective preference lists). Due to the shortening of the preference lists as the algorithm progresses (see Algorithm˜2), this is referred to as the version of the DA algorithm with Shortlists [ESM]. While this version terminates with at most as many (and often less) proposal steps as compared to Algorithm˜1, it requires repeated updates to the original preference lists.
@@
C.2 Algorithm for Generating a Stable Matching w/ Master-Lists
Similar to Algorithm˜1, Algorithm˜3 runs in rounds of proposals. Since there is a Master-list over proposing agents in the preference lists of receiving agents, and proposing agents are selected to make proposals in the order in which they appear in the Master-list, there are no rejections (since any receiver receives the best possible proposal at any step) [irving2008stable]. Hence, this algorithm terminates in $n$ proposal steps (and is therefore easier to execute).
Algorithm 3 The Deferred-Acceptance Algorithm for Preferences w/ Master-lists on One Side
@@ C.3 Algorithm for Resolving Instability
While we don’t explicitly describe the algorithm here, the mechanism presented by abeledo1995paths can be applied to an unstable matching $\mu$ by resolving blocking pairs, resulting in a stable solution. Intuitively, an LLM does not have to follow a specific mechanism, rather the model can resolve instability by iteratively resolving blocking pairs as they arise (eventually, assuming all steps are correct, the model should arrive at a stable solution).
@@ C.4 Algorithm for Detecting Instability
Intuitively, Algorithm˜4 works by iteratively visiting each pair of agents $(m,w)$ s.t. $m\in M$ and $w\in W$, and finding a pair such that either $m$ prefers $w$ to their current partner, or $w$ prefers $m$ to their current partner (when such a pair is found, output it as a blocking pair). If no pair $(m,w)$ is found to be a blocking pair, then the solution is stable.
@@ Appendix D Prompt Engineering
In Table˜4, examine LLMs’ performance when provided with a prompt containing pseudocode for the DA algorithm, and compare it to the case when no algorithm is provided in the prompt. While providing the DA algorithm as part of the prompt leads to an improvement in some cases, the only case in which there is improvement is statistically significant [^8] At $p<0.05$, using Fisher’s exact test is with DeepSeek-70B at the Easy level with ML instances.
@@
For models that fail at generating stable matchings even with small instances, we evaluate whether prompt-based enhancements such as Chain-of-Thought (CoT) [wei2022chain] and Few-shot prompting [dong2022survey] can improve their performance. In particular, we introduce the following three types of modifications to the original prompt (used in Section˜3):
  • CoT-Vanilla (CoT-V): The steps of execution of the DA algorithm (see Algorithm˜1) are provided for an example instance. Each step consists of a single (free) proposing agent making his next proposal, and the receiving agent either accepting or rejecting the proposal based on their current status.
  • CoT-Shortlist (CoT-SL): This version of CoT uses the Shortlist version of the DA algorithm (see Algorithm˜2) which reduces the overall number of proposal steps, but requires repeated updates to the original preference lists.
@@ 20 IC 0 0 0 0 0 0 0 0 14 2 6 6 0 0 0 0
As depicted in Table˜5, these prompting enhancements fail to improve the ability of LLMs to generate stable matchings. While there is a slight improvement for models like Qwen-QwQ-32B and DeepSeek-70B instances with Master-list preferences and size $n=10$, this improvement is not statistically significant.
@@
We consider the traditionally used setting of stable-marriage, where the set $M$ consists of men who propose to women in the set $W$ [GS62college]. To measure whether LLMs are sensitive to the nomenclature used to described the two-sided matching market, we also consider a different setting, i.e. that where a set of workers ($W$) are to be assigned a set of tasks ($T$) (and members on both sides have preferences over members of the other). We test the difference in the performance of two LLMs—Gemini-2.0-Flash and o3-mini—between the task-scheduling and stable-marriage settings. The results are provided in Table˜6. While there is a slight decrease in performance for o3-mini, the change is not significant (at $p<0.05$). This provides further evidence that LLMs understand requirements of computing stable solutions for matching markets, in general.
@@ 50 IC 0 0 80 72
Appendix E Generating Stable Solutions for Preferences with Ties
As demonstrated in Section˜3, reasoning-enabled models achieve significantly higher accuracy on Easy instances when compared to non-reasoning baseline models, but suffer dramatic drops in performance with Hard instances. A natural extension of this is to examine how introducing ties to preferences impacts an LLM’s ability to generate stable solutions.
New notions of stability.
When ties are introduced to preference profiles, additional (stronger) notions of stability exist. Namely, in addition to the standard notion of weak stability (where a matching does not admit any weak blocking pairs: agents who strictly prefer each other to their current partners), we have the added notions of strong and super stability, where a matching does not admit any strong or super blocking pairs, respectively. A strong blocking pair is a pair of agents in which one agent strictly prefers the other agent over their current partner, and the other agent remains indifferent or prefers the first agent and their current partner. A super blocking pair is a pair of agents in which either agent is either indifferent the other agent and their current partner, or prefers the other agent to their current partner. Super stability, the strongest notion of stability, implies strong stability, which in turn implies weak stability. Additionally, it is important to note that strongly and super stable solutions do not exist for all preference profiles.
Irving [IRVING1994261] provides three algorithms (one for each) to compute weakly, strongly, and super stable matchings for preference profiles with ties. The algorithm for weak stability takes $\mathcal{O}(n^{2})$ steps (the algorithm is equivalent to running standard the DA algorithm with arbitrary tie-breaking), the algorithm for strong stability takes $\mathcal{O}(n^{4})$ steps, and the algorithm for super stability takes $\mathcal{O}(n^{2})$ steps (these two algorithms require more demanding steps with complex operations).
Figure 7: The generated responses by Gemini-2.5-Pro for Master-list (ML) and Impartial Culture (IC) Easy preferences with different prompts. All generated matchings were valid and complete. Again, unstable indicates one-to-one matchings with weak blocking pairs present.
Generating matchings for each stability notion.
To evaluate how LLMs handle preferences with ties, we perform the same experiment as in Section˜3 with Easy preference profiles and Gemini-2.5-Pro, except we introduce multiple ties of random sizes, at randomly selected starting positions, in each preference list in the profile. Additionally, we modify the prompt listed in Section˜J.1 to get a total of three different prompts. The first prompt (referred to as the Baseline Prompt) is identical to the prompt in Section˜J.1. The other two prompts (referred to as the Strong Prompt and Super Prompt) replace the instruction to “…find the proposer-optimal stable matching…” with “…find the proposer-optimal STRONG stable matching…” and “…find the proposer-optimal SUPER stable matching…”, respectively. For each prompt, we run the experiment and count the proportion of generated matchings that are super, strong, and weakly stable, as well as unstable matchings.
Figure˜7 demonstrates the performance of Gemini-2.5-Pro in generating stable outcomes with respect to each of the three new notions of stability for preference profiles with ties. Gemini-2.5-Pro once again demonstrates a high accuracy ($96\%$) in generating stable solutions (considering weak, strong, and super stability) with the Baseline Prompt, with $50\%$ and $54\%$ of stable solutions being weakly stable with the IC-Easy and ML-Easy preference profiles, respectively. However, only around $40\%$ of generated solutions were super (and strongly) stable for both preference cultures. Gemini-2.5-Pro’s reasoning traces for each instance indicate that the model essentially attempts to execute the DA algorithm when generating stable solutions even when ties are present (rather than utilizing the more complex algorithms for strong and super stability, even when explicitly prompted). Additionally, any generated super stable solutions are a result of an instance’s super stable solution intersecting with its weakly stable solution. Interestingly, specifying the notion of stability appears to significantly hurt the model’s ability to generate any stable solutions. While one potential explanation is that specifying the desired stability notion in the prompt introduces unnecessary noise, this is an interesting avenue for future study.
Appendix F Resolving Instability: Additional Material
F.1 Generating Unstable Matchings
@@
We generate a matching that contains a single blocking pair, by swapping the partners of two randomly selected proposer agents in the Optimal matching. Since such a swap may lead to more than one blocking-pair (or no blocking pairs), we perform this process (for every instance) until we obtain a matching with exactly one blocking-pair. This procedure is formally described in Algorithm˜5.
@@
[…]
@@
[…]
8:   $\mu\leftarrow\text{copy}(\mu^{\ast})$ $\triangleright$ Start from the stable matching
@@
11:   $w_{b}\leftarrow\mu(m_{b})$ $\triangleright$ Swap partners of the two men
@@
F.2 Further Results
Table 7: Instability Rate (averaged across instances) in the (valid) matchings returned by LLMs when asked to resolve a given unstable matching of types One-BP and Random. The column “Baseline” contains the (average) Instability Rate for the provided matching of the indicated type. Numbers in bold indicate that the Instability Rate of the corrected solution is significantly worse than the provided matching. A * on the number in the One-BP column indicates a that Instability Rate is significantly lower than the case when a Random matching is provided (at p<0.05p<0.05). [^12] All statistical comparisons in this table are made using Welch’s t-test [WelchTTest].
@@ F.2 Further Results
The extent to which a given matching is incorrect does influence quality of the solutions returned by LLMs after being asked to correct it. Table˜7 shows that the number of blocking pairs is often significantly lower when the unstable matching (that LLMs are asked to correct) has a single blocking pair, as compared to the case with random matchings. [^13] In fact, asking LLMs to resolve a random matching leads to a significantly higher number of blocking pairs in the returned solution, as compared to the case when they are asked to generate solutions from scratch. Similarly, as shown in Figure˜8, matchings returned after resolving an almost stable matching have a greater overlap with the Optimal solution (especially at Medium and Hard difficulty levels).
Appendix G Additional Material about Detecting Instability
G.1 Comparing Impartial Culture and Master-List Instances
@@
Appendix H Fine-tuning Details
@@
We fine-tune four reasoning models:
@@
  • DeepSeek-14B (deepseek-ai/DeepSeek-R1-Distill-Qwen-14B),
  • Qwen-QwQ-32B (Qwen/QwQ-32B), and
  • Qwen3-32B (Qwen/Qwen3-32B)
using the Unsloth [^14] https://unsloth.ai/ framework with parameter-efficient tuning (LoRA). [^15] Since results for Qwen-QwQ-32B and Qwen3-32B are similar, we show only those for the former in Table 2.
@@
The dataset for the Generation task contains $N=10000$ samples for which the reasoning trace is generated using a Python implementation of the DA algorithm. For the Preference Reasoning task, the dataset consists of $N=9000$ samples ($3000$ for each question level), where the reasoning trace involves explicitly identifying the positions of agents in the concerned preferences. In both datasets, we include an equal number of IC and ML instances, with sizes ranging from $n=5$ to $n=50$. Detailed examples of training examples for both tasks (Generation and Preference Reasoning) are provided in Appendix˜K.
@@
Ablation tests.
Table˜8 illustrates how performance is sensitive to variations in data-related parameters such as the types of instances included in the data, range of instance sizes, and number of training examples, as well as training-related parameters such as the LoRA rank and the base model used. The primary observation from these tests is that easier training instances are more crucial that harder ones. For example, performance is significantly better when the training data consists of only Easy (and smaller) instances, i.e. with $n=10$, as compared to the case with only Hard instances ($n=50$). Similarly, performance is much better when the data consists of only ML instances, as opposed to the case where it consists of only IC instances, which are more difficult than the former.
Table 8: Performance (percentage of stable solutions generated) after fine-tuning, for DeepSeek-8B and Qwen3-32B, with different configurations of data- and training-related hyper-parameters. Configuration 1 is the default configuration. The variation in each other configuration (as compared to config. 1) is in bold.
ML IC
Config. Instance types Instance sizes (range) Training set size LoRA rank Base model Easy Medium Hard Easy Medium Hard Total
1 ML, IC [5,50] 10000 32 DeepSeek-8B 94 60 0 34 2 0 31.67
Qwen3-32B 100 98 4 98 84 0 64
2 ML [5,50] 10000 32 DeepSeek-8B 100 94 0 2 0 0 32.67
Qwen3-32B 100 100 0 6 0 0 34.33
3 IC [5,50] 10000 32 DeepSeek-8B 0 0 0 0 0 0 0
Qwen3-32B 0 0 0 0 2 0 0.33
4 ML, IC [5,50] 10000 64 DeepSeek-8B 96 82 0 2 0 0 30
Qwen3-32B 100 100 0 100 100 0 66.67
5 ML, IC [5,50] 5000 32 DeepSeek-8B 74 46 0 0 0 0 20
Qwen3-32B 100 94 0 98 92 0 64
6 ML, IC [5,10] 5000 32 DeepSeek-8B 100 2 0 100 8 0 35
Qwen3-32B 100 52 0 100 26 0 46.33
7 ML, IC [50,50] 5000 32 DeepSeek-8B 0 0 0 0 0 0 0
Qwen3-32B 12 0 0 2 0 0 2.33
Appendix I Inference Details
@@
Appendix J Prompts
J.1 Example prompt for Generating Stable Solutions
@@
The prompt used for generating stable solutions with LLMs follows standard prompting procedures, by first outlining the task, providing appropriate context, specifying constraints, and detailing the desired output format (a JSON object). We intentionally provide preferences in a structured, tabular format. It enables us to isolate and rigorously evaluate the LLMs’ reasoning, alignment, and solution quality relative to normative axioms (e.g., stability). Natural language formulations introduce significant noise in both input and output, making it difficult to attribute performance failures to reasoning versus parsing. By grounding our analysis in tabular settings first, we can obtain clean and interpretable measurements, forming a benchmark for future extensions that incorporate naturalistic input.
Notice that despite the deferred-acceptance algorithm never being mentioned in the prompt, all models mentioned the deferred-acceptance algorithm in their responses. As mentioned in Appendix˜D, we use the traditional setting of stable-marriage (where men propose to women) considered by gale1962college while describing the problem in the prompt.
You are an intelligent assistant who is an expert in algorithms. Consider the following instance of the two-sided matching problem, where 10 men are to be matched with 10 women. Here are the preference lists for all individuals: <<preferences>> { M: { M1: [W10,W1,W3,W6,W2,W4,W9,W8,W7,W5], M2: [W8,W3,W10,W6,W2,W5,W4,W7,W1,W9], M10: [W2,W5,W1,W3,W7,W6,W10,W4,W9,W8], }, W: { W1: [M2,M8,M9,M10,M5,M7,M1,M4,M6,M3], W2: [M2,M7,M3,M1,M8,M9,M6,M10,M5,M4], W10: [M6,M4,M7,M5,M8,M9,M10,M2,M3,M1], }} <</preferences>> Your task is to find the proposer-optimal stable matching. You can use XML tags like <<scratchpad>> to explain your thought process while computing the solution. Once you have found a stable matching, please return your matching in the JSON format given below: <<answer>> { “M1”: “<<woman matched with M1>>”, “M2”: “<<woman matched with M2>>”, “M10”: “<<woman matched with M10>> } <</answer>> Make sure that each man/woman is matched with exactly ONE partner. It is mandatory that you provide a matching as a JSON object enclosed in <<answer>><</answer>> tags as described above.
@@
You are an intelligent assistant who is an expert in algorithms. <</preferences>> Your task is to find the proposer-optimal stable matching. For this, you can use the Deferred Acceptance algorithm. The steps of this algorithm are described below: 1. Initialize all men and women as unmatched. 2. Create a list to keep track of each man’s next proposal (initially set to 0 for all men). 3. While there are unmatched men: a. Select an unmatched man (M). b. Find the next woman (W) on M’s preference list that he hasn’t proposed to yet. c. If W is unmatched, match M and W. d. If W is matched but prefers M to her current partner: - Unmatch W from her current partner. - Match M and W. - Set the unmatched man as W’s previous partner. e. If W rejects M, move to the next woman on M’s preference list. 4. Repeat step 3 until all men are matched. You can use XML tags like <<scratchpad>> to explain your thought process … It is mandatory that you provide a matching as a JSON object enclosed in <<answer>><</answer>> tags as described above.
@@
You are an intelligent assistant who is an expert in algorithms. Consider the following instance of the two-sided matching problem, where 5 workers are to be assigned with 5 tasks, and each worker is assigned exactly one task. Here are the preference lists for all workers (W) over tasks (T) and the preferences of tasks over workers: <<preferences>> { W: { W1: [T5, T3, T4, T2, T1] } T: { T1: [W3, W5, W4, W1, W2] }} <</preferences>> Your task is to find a stable matching of workers and tasks. You can use XML tags like <<scratchpad>> to explain your thought process while computing the solution. Once you have found a stable matching, please return your matching in the JSON format given below: <<answer>> { "W1": "<<task assigned to W1>>", "W5": "<<task assigned to W5>>" } <</answer>> Make sure that each worker is assigned exactly ONE task. It is mandatory that you provide a matching as a JSON object enclosed in <<answer>><</answer>> tags as described above.
J.2 Example Prompts for Prompt Engineering
J.2.1 CoT-Vanilla
@@
Chain-of-Thought methods were applied to the prompt in Section˜J.1 by additionally including an example trace of steps performed when running the deferred-acceptance algorithm on a randomly generated instance. The algorithm trace includes all proposals, all respective acceptances/rejections, and the resultant stable solution. The entire Chain-of-Thought example is enclosed within <<example>> XML tags.
You are an intelligent assistant who is an expert in algorithms. Your task is to find the proposer-optimal stable matching, for the two-sided matching problem. Here is an example to demonstrate how you should proceed: <<example>> <<preferences>> { M: { M1: [W5,W1,W2,W4,W3], M2: [W1,W2,W5,W4,W3], M3: [W4,W2,W3,W1,W5], M4: [W5,W1,W2,W4,W3], M5: [W3,W5,W4,W2,W1], }, W: { W1: [M2,M3,M5,M4,M1], W2: [M5,M2,M4,M3,M1], W3: [M2,M1,M3,M5,M4], W4: [M1,M4,M5,M3,M2], W5: [M4,M3,M5,M2,M1], }} <</preferences>> M1 is free. M1 proposes to W5 Since W5 is free, W5 accepts the proposal. Now M1 and W5 are matched. M2 is free. M2 proposes to W1 Since W1 is free, W1 accepts the proposal. Now M2 and W1 are matched. M3 is free. M3 proposes to W4 Since W4 is free, W4 accepts the proposal. Now M3 and W4 are matched. M4 is free. M4 proposes to W5 Since W5 prefers M4 to their current partner M1, W5 accepts the proposal. Now M4 and W5 are matched, and M1 is free. M1 is free. M1 proposes to W1 Since W1 prefers their current partner M2 to M1, W1 rejects the proposal. M2 and W1 are still matched, and M1 is still free. M1 is free. M1 proposes to W2 Since W2 is free, W2 accepts the proposal. Now M1 and W2 are matched. M5 is free. M5 proposes to W3 Since W3 is free, W3 accepts the proposal. Now M5 and W3 are matched. <<answer>> { "M1": "W2", "M2": "W1", "M3": "W4", "M4": "W5", "M5": "W3" } <</answer>> <</example>> Consider the following instance of the two-sided matching problem, where 10 men are to be matched with 10 women $\ldots$
J.2.2 CoT-Shortlist